Two parallel large thin metal sheets have equal surface charge densities of opposite signs. The electric field between these sheets is..
( charge per unit length= 26.4 * 10raised to the power -12 c/m2...
Please provide answer with solution

Grade:12th pass

1 Answers

Arun
5 years ago
Electric field intensity between two parallel sheets have equal surface charge densities of opposite sign is given by E = σ/ε₀
Here given, σ = 26.4 × 10⁻¹²C/m² 
and ε₀ = 8.875 × 10⁻¹² C²/Nm² 
So, Electric field intensity between sheets = 26.4 × 10⁻¹²/8.875 × 10⁻¹² ≈ 3 N/C 
Hence, answer is 3N/C